OpenStudy (anonymous):

Find the derivative of the function h(x) = 8^x(squared) -x

OpenStudy (anonymous):

\[8^{x^2-x}\]

OpenStudy (anonymous):

chain rule, plus knowing \(\frac{d}{dx}b^x=b^x\ln(b)\)

OpenStudy (anonymous):

how would i start the problem though?